### v2

# Event-based Sensor Fusion and Application on Odometry: A Survey

Abstract
Event cameras, inspired by biological vision, are asynchronous sensors that detect changes in brightness, offering notable advantages in environments characterized by high-speed motion, low lighting, or wide dynamic range. These distinctive properties render event cameras particularly effective for sensor fusion in robotics and computer vision, especially in enhancing traditional visual or LiDAR-inertial odometry. Conventional frame-based cameras suffer from limitations such as motion blur and drift, which can be mitigated by the continuous, low-latency data provided by event cameras. Similarly, LiDAR-based odometry encounters challenges related to the loss of geometric information in environments such as corridors. To address these limitations, unlike the existing event camera-related surveys, this paper presents a comprehensive overview of recent advancements in event-based sensor fusion for odometry applications particularly, investigating fusion strategies that incorporate frame-based cameras, inertial measurement units (IMUs), and LiDAR. The survey critically assesses the contributions of these fusion methods to improving odometry performance in complex environments, while highlighting key applications, and discussing the strengths, limitations, and unresolved challenges. Additionally, it offers insights into potential future research directions to advance event-based sensor fusion for next-generation odometry applications.
Index Terms:
Event camera, Sensor fusion, Odometry, Visual Sensors, LiDAR
I Introduction
Event cameras are bio-inspired vision sensors that operate fundamentally differently from traditional frame-based cameras. Instead of capturing images at fixed time intervals, event cameras detect changes in brightness asynchronously, resulting in a stream of events that encode changes in the scene [1, 2]. This unique sensing modality provides significant advantages in scenarios involving high-speed motion, low lighting, or scenes with significant dynamic range, where traditional cameras may struggle [3].
In the context of robotics and computer vision, sensor fusion involving event cameras has garnered considerable attention, particularly for visual odometry (VO) or visual-inertial odometry (VIO). Visual odometry aims to estimate a robot’s trajectory by analyzing visual inputs, often combined with other sensor data. Event-based sensor fusion, therefore, seeks to leverage the complementary strengths of event cameras and other sensors (e.g., frame-based cameras, inertial measurement units (IMUs), LiDAR) to improve robustness and accuracy in hash environments with high-speed needs. Regarding odometry, LiDAR-based systems generally offer greater accuracy and robustness to varying lighting conditions compared to visual odometry (VO) or visual-inertial odometry (VIO). However, in environments where geometric information is homogenous, such as long corridors, LiDAR odometry is prone to drift [4]. Event cameras, with their ability to capture dynamic scene changes, have the potential to mitigate this issue by providing complementary information that enhances the overall accuracy and robustness of odometry in such challenging scenarios. Figure 1 illustrates the varying representations of a long corridor as captured by different sensors, including an RGB camera, LiDAR, and an event camera. It is evident from the LiDAR data that the corridor is not clearly detected; however, by integrating the data from the RGB camera and the event camera, this limitation can be mitigated.
Figure 1: The sensor data representation of a long corridor includes the following: RGB image (top-left), an event-based camera image (middle-left), a LiDAR point cloud (top-right), and a LiDAR-generate reflectivity image (bottom). LiDAR data is adapted from [5].
The primary motivation for using event cameras in sensor fusion is their ability to provide continuous, low-latency information that complements the limitations of traditional sensors. For instance, frame-based cameras are prone to motion blur during rapid movements, while IMUs suffer from drift over time. By fusing event camera data with other modalities, it is possible to enhance the accuracy of visual odometry, particularly in challenging environments such as fast-changing scenes or low-light conditions.
Unlike other event-camera-based surveys [3, 6, 7, 8, 9], we are particularly interested in the integration of even cameras for odometry purposes. This survey aims to provide a most comprehensive and systematic overview of recent advancements in event-based sensor fusion and its application to visual odometry. We explore different fusion strategies, including event and frame-based camera fusion, event and IMU fusion, and event and LiDAR fusion, examining their contributions to improving visual odometry performance. We also discuss various methodologies, highlight important applications, and analyze the strengths and limitations of these approaches. Additionally, we identify open challenges and promising directions for future research, providing insights into how event-based sensor fusion can be further developed to meet the needs of next-generation robotic systems.
The remainder of this paper is structured as follows: Section II provides an overview of the background and fundamentals of event cameras. Following that, Section III discusses the key trends in event camera-based sensor fusion tasks particularly on enhancing odometry by integrating event cameras with other sensors. Section IV introduces the most recent datasets relevant to event cameras for odometry. Finally, Section V concludes the paper, offering insights into potential future research directions.
Figure 2: Primary aspects covered in event camera-based fusion for odometry purposes in this survey.
II Background of the Event Camera
Event cameras capture changes in brightness asynchronously for each pixel. Unlike traditional frame-based cameras that capture intensity images at fixed intervals, event cameras detect changes in the logarithmic brightness of individual pixels independently, producing events whenever a change exceeds a predefined threshold [6]. Specifically, an event at a pixel is triggered if the log intensity changes since the last event surpasses the contrast threshold:
$\Delta L(x,y,t)=\log I(x,y,t)-\log I(x,y,t_{\text{last}})=p\cdot C$ (1)
where $\Delta L(x,y,t)$ represents the logarithmic brightness at the pixel at a time; $I(x,y,t)$ is the intensity at pixel $(x,y)$ at time $t$; $t_{\text{last}}$ is the time of the last event at pixel $(x,y)$; $p\in\{+1,-1\}$ represents the polarity of the event; $C$ is the contrast threshold. This threshold determines the camera’s sensitivity to brightness changes.
Currently, there are multiple primary event camera manufacturers which are listed in Table I alongside several key features illustrated.
TABLE I: Event Camera Products
Brand Main Products Key Features Stereo Kit
Prophesee EVK4, EVK3 High Resolution
IniVation DAVIS346, DVXplorer Larger community
Lucid Triton2 EVS Industrial application, PoE, GigE
Sony IMX636, IMX637, IMX646, IMX647 Sensor manufacturer
II-A Comparison with Traditional Cameras
Event cameras offer several advantages over traditional frame-based cameras, which make them particularly suitable for specific applications in robotics and computer vision [3, 6]. Unlike traditional cameras that capture images at fixed frame rates, event cameras capture changes in brightness asynchronously for each pixel. This difference leads to several key advantages:
  • High Temporal Resolution: Event cameras only generate data when a change occurs in the scene, resulting in sparse, storage-efficient output, unlike traditional cameras that produce dense frames at fixed intervals regardless of scene changes.
  • Low Latency: Each pixel in an event camera operates independently, transmitting brightness changes immediately without global exposure, resulting in minimal latency (10 ms in lab tests and sub-milliseconds in real-world conditions).
  • High Dynamic Range: With a dynamic range of up to 140 dB (compared to 60-70 dB in traditional cameras), event cameras perform well in environments with significant lighting variations, avoiding saturation and underexposure.
  • Reduced Motion Blur: Event cameras capture asynchronous changes at each pixel, making them inherently resistant to motion blur, especially in fast-moving or low-light conditions, ideal for visual odometry in dynamic environments.
Despite these advantages, event cameras also present challenges, such as the difficulty of interpreting event streams and the need for specialized algorithms to process asynchronous data. Traditional frame-based cameras, on the other hand, provide rich spatial information that is easier to use for tasks like feature extraction and object recognition, which makes them more suitable for certain applications where high spatial resolution is required.
The complementary nature of event cameras and traditional cameras has led to the development of fusion approaches that leverage the strengths of both modalities. By combining the high temporal resolution and dynamic range of event cameras with the structured spatial information from traditional cameras, it is possible to achieve more robust and accurate perception systems.
II-B Sensor Fusion for Odometry in Robotics
Sensor fusion involves integrating data from multiple sensors to achieve more reliable and accurate measurements. In the context of odometry estimation for robotics and autonomous systems, several key sensors are commonly utilized, including visual sensors, LiDAR, IMU, and others.
Traditional VO methods can suffer from challenges such as motion blur, occlusion, and sensitivity to lighting variations. Event cameras, with their high temporal resolution and ability to capture asynchronous changes, offer a promising solution to address these challenges by providing complementary information that enhances the robustness and accuracy of VO.
To enhance robustness, several studies have explored different sensors like LiDAR, IMU, and radar. Fast-LIO [10][11] combined LiDAR and IMU sensors to achieve odometry, where LiDAR points were extracted as planar and edge features. An iterative Kalman filter was used to fuse these features with IMU measurements, providing accurate state estimation. Building on this, COIN-LIO [12] introduced an improvement by incorporating photometric information from LiDAR intensity data. The system projected LiDAR intensity values into an image format and applied filtering to enhance brightness consistency. The key features in these intensity images were selected based on their complementarity to the geometric features of the LiDAR point cloud, specifically targeting areas with limited geometric information.
Camera images, on the other hand, offer rich semantic information, making them crucial for robotic perception. To leverage this, camera data is often fused with other sensors such as IMUs (providing acceleration and angular velocity) or LiDAR (providing precise depth information). This fusion significantly improves odometry performance, particularly in complex environments or under poor lighting conditions. Frameworks like VINS-Mono [13], which fuses camera and IMU data, and Kimera-VIO [14, 15], which integrates LiDAR, camera, and IMU data, exemplify the advantages of such approaches. These systems enable more reliable navigation, even in low-texture or poorly lit environments, resulting in more accurate and robust odometry for real-world applications.
III Event-based Sensor Fusion for Odometry
III-A Event-Camera Only
Since the Event camera only detects the change in brightness, the features that can be used for localization are less than frame cameras. Hence, the amount of studies using only an event camera for odometry is small. The first event-based pose estimation work was presented by Weikersdorfer et al. [16] in 2012. This work used only an event camera and was extended to an event-based visual SLAM system in 2013 [17].
III-B Fusion with Frame-based Camera
Combining event cameras with frame-based cameras allows for high temporal resolution with more structured scene information. In [18], the authors present an approach to combine both sensor modalities for object detection and tracking. In [19], the authors propose a method to reconstruct high frame-rate, sharp videos from a single blurry frame and event data using the Event-based Double Integral (EDI) model. The fusion of asynchronous event data with frame-based data reduces blur and improves visual quality, showcasing the complementary strengths of both modalities. Fusing an event camera with a frame camera can overcome the limitation of each separate sensor and estimate the stereo depth. For example, the work in [20] enables stereo depth estimation by leveraging both event-based edge information and frame-based intensity data, offering improved disparity estimation compared to state-of-the-art stereo matching algorithms.
The fusion of an event camera and a frame camera can increase the robustness and accuracy of visual odometry (VO). In 2014, Andrea Censi and Davide Scaramuzza proposed a low-latency event-based VO system [21], which was a significant advancement in reducing the latency associated with traditional frame-based methods. This system utilized asynchronous event data to compute odometry in real time, paving the way for faster and more efficient pose estimation in dynamic environments. Then in 2016, the work further enhanced the field by introducing low-latency VO using event-based feature tracks [22]. This method tracked feature points based on event data, achieving reliable and low-latency odometry even in conditions where traditional frame-based methods would struggle, such as in fast motion or poor lighting. The same group developed a 6-degree-of-freedom (DoF) event-based tracking and mapping system, which relied on geometric principles to perform both parallel tracking and mapping in real time [23]. This work was instrumental in showcasing the potential of event cameras for performing full 3D tracking and mapping tasks, marking another milestone in event-based VO.
In 2022, Hidalgo-Carrión et al. proposed event-aided direct sparse odometry (DSO), where event data was combined with traditional DSO to improve the robustness of VO in environments with challenging lighting or rapid motion changes [24]. This method demonstrated the fusion of event-based and frame-based approaches in improving VO performance.
III-C Fusion with IMU
IMUs provide inertial data that can be fused with events to enhance motion estimation, particularly in high-speed scenarios. IMUs can provide estimations of velocity and orientation, aiding visual odometry by reducing drift and improving robustness [25]. In [26], the authors explore the use of event data and IMU fusion to achieve enhanced VO performance in dynamic environments. The study proposes methods to leverage the event camera’s high temporal resolution with the IMU’s robust inertial data to accurately estimate motion, even in highly dynamic scenarios Another notable work is [27], the authors propose new methods for integrating event data and IMU readings to achieve ultra-fast camera pose estimates. They introduce two novel fusion schemes that combine constant velocity and constant acceleration prediction models with ultra-fast (10 kHz) event-based updates and slower 1 kHz IMU updates. The findings demonstrate superior performance at a throughput that is 100 times faster than the state-of-the-art.
Integrating an IMU into VO systems significantly enhances their performance by providing additional, reliable motion information that compensates for the limitations of visual data alone. In 2017, Zhu et al. extended the capabilities of event-based odometry by integrating inertial measurements in their work on event-based visual-inertial odometry (EVIO) [28]. The combination of event cameras and inertial sensors allowed for robust pose estimation even in challenging scenarios, such as high-speed motion or complex dynamic scenes. In 2018, Elias Mueggler and his team presented a continuous-time visual-inertial odometry approach for event cameras, further enhancing the accuracy and robustness of event-based odometry systems [29]. Their work highlighted the importance of fusing temporal information from both event data and inertial sensors in a continuous-time framework.
Florian Mahlknecht et al. explored event camera-based odometry for planetary robots in 2022, showcasing the potential of event cameras in extraterrestrial environments where traditional cameras may struggle due to harsh conditions [30]. In 2023, the authors in [31] introduced a hybrid approach combining event and frame-based visual-inertial odometry with adaptive filtering techniques based on 8-DoF warping uncertainty. This method enhanced the odometry system’s resilience to noise and dynamic changes, further improving the robustness and adaptability of event-based systems in various application settings.
III-D Event-based Stereo Visual Odometry
Event-based stereo visual odometry is an evolving field that leverages the advantages of stereo event cameras, which provide asynchronous data from two synchronized event sensors, allowing for 3D depth perception in challenging environments with high-speed motion or varying lighting conditions. The early works on Event-based stereo VO were published in 2018. Henri Rebecq et al. [32] introduced EMVS (Event-based Multi-View Stereo), which enabled 3D reconstruction in real-time using event cameras. This method capitalized on the asynchronous nature of events to compute depth information by fusing event data from multiple viewpoints, enabling accurate and efficient 3D reconstruction. In the same year, a semi-dense 3D reconstruction method using stereo event cameras was proposed [33]. Their approach utilized stereo event cameras to reconstruct semi-dense 3D maps, significantly improving the accuracy and robustness of visual odometry systems, even in dynamic environments. In [34], the authors presented the Multivehicle Stereo Event Camera Dataset, specifically designed for 3D perception. This dataset provided benchmark data for stereo event-based odometry, enabling further research in event-based 3D perception. By providing a large-scale, real-world dataset, this work facilitated the evaluation of stereo event camera systems in various challenging scenarios.
In 2021, Zhou et al. [35] extended their prior work in [33] by introducing a complete event-based stereo visual odometry system. Their method combined stereo event cameras and a robust odometry pipeline to estimate camera motion in real time. By leveraging the high temporal resolution of event cameras, this system was able to track fast-moving objects and scenes with challenging lighting conditions. The feature-based method can also be utilized on event-based stereo visual odometry [36]. Their approach detected and tracked features directly from stereo event data, using them to estimate motion.
Recently, research in stereo event-based visual-inertial odometry (VIO) has gained attention, as it enhances motion estimation by fusing inertial measurements with visual data to improve accuracy and robustness, particularly in conditions where visual information alone may be unreliable. In 2023, Kunfeng Wang, Kaichun Zhao, and Zheng You developed a stereo event-based VIO system, which incorporated both stereo event cameras and inertial measurements to improve the accuracy of pose estimation [37]. Further research into stereo event-based VIO led to the work of Peiyu Chen et al. [38]. In parallel, Liu et al. also presented an event-based stereo visual-inertial odometry system [39]. Their methods demonstrated superior performance in both accuracy and robustness when compared to frame-based VIO systems, particularly in scenarios involving rapid motion.
Most recent works by Junkai Niu et al. have continued to push the boundaries of event-based stereo VO. Their IMU-aided event-based stereo visual odometry system[40] and ESVO2 [41] incorporate direct visual-inertial odometry with stereo event cameras, significantly improving real-time pose estimation accuracy.
TABLE II: Event camera-based dataset related to odometry
Year Dataset Name Scenarios Sensors
2024 CoSEC [42] Driving scenarios Event Camera, LiDAR, RGB Camera, IMU, GNSS
2023 ECMD [43] Driving scenarios Event Camera, LiDAR, RGB Camera, Infrared Camera, IMU, GNSS
2022 Evimo2 [44] Indoor Scenes Stereo Event Camera, IMU
2021 DSEC [45] Driving scenarios Stereo Event Camera, LiDAR, RGB Camera, IMU, GNSS
2017 DAVIS Datasets [46] High-speed robotics Event Camera, RGB Camera, IMU
III-E Fusion with LiDARs
LiDAR technology generates precise 3D maps and is capable of producing highly accurate odometry information. However, it operates at a lower frequency and encounters challenges in specific scenarios, as previously mentioned, where geometric information is compromised or lost. Several studies have focused on the fusion of event cameras with LiDAR systems, aiming to leverage the advantages of event cameras to enhance LiDAR performance.
Some of the research is still in the early stages, with a primary focus on achieving automatic and accurate calibration. In [47], the authors introduce a direct extrinsic calibration method between event cameras and LiDAR, exploiting the high temporal resolution and dynamic range of event cameras. This technique eliminates the need for frame-based intermediaries and optimizes 6-DoF extrinsic calibration through information-based correlation methods, improving sensor alignment. Similarly, in [48], a novel calibration method is proposed, using a 3D marker to generate stable patterns detectable in both 2D point sets and 3D point clouds. This approach facilitates accurate calibration between the event camera and LiDAR by leveraging light changes to trigger event detection, which is crucial for achieving stable fusion. This paves the way for the fusion of event cameras and LiDARs.
Beyond calibration, LiDAR and event camera fusion can enhance depth perception. In [49], a method proposes fusing sparse LiDAR and event data to generate dense depth maps, improving spatial perception. This improved point cloud map can further facilitate the LiDAR-based odometry. In the work [50], the authors discovered that data from event cameras share a homogeneous nature with RGB and LiDAR data in both visual and motion spaces. This characteristic makes event cameras particularly suitable for bridging the gap between RGB and LiDAR in motion fusion or scene flow tasks. The finding highlights the potential for a hierarchical fusion framework that effectively integrates these sensors, making the fusion of event cameras with other sensors like LiDAR and conventional cameras highly promising.
IV Event Camera Dataset
Unlike other reviews of event camera-based datasets, this part of the paper aims to highlight several event camera-related datasets that have particularly been developed to support research in event-based sensor fusion for odometry, enabling accurate pose estimation and motion tracking in dynamic environments. The datasets, along with the sensors involved, are listed in Table II.
V Conclusion and Discussion
V-A Conclusion
Due to their high temporal resolution, low latency, high dynamic range, and reduced motion blur compared to traditional cameras, event cameras offer significant potential benefits for odometry calculation by capturing precise changes in brightness. Unlike previous review papers focused on event cameras, this paper provides a comprehensive and detailed analysis of the application of event cameras for odometry, exploring various sensor fusion approaches. These include event camera-only systems, fusion with IMUs, stereo event camera setups, integration with traditional cameras, and combinations with LiDAR. Additionally, we briefly discuss event camera-related datasets that are relevant for advancing sensor fusion, particularly in odometry applications.
V-B Discussion
Event camera holds significant potential for future applications across various domains. Its role in odometry or pose estimation has been specifically addressed in the paper. In Visual Odometry (VO) or Visual-Inertial Odometry (VIO), the use of event cameras can effectively mitigate motion blur and reduce the challenges posed by low-light environments, thanks to their wider dynamic range. For LiDAR-based odometry, similar to their application with conventional cameras, event cameras can help reduce drift by compensating for the loss of geometric information in LiDAR point clouds with photometric information. Additionally, event cameras can be integrated with LiDAR-generated images, as modern LiDAR systems often produce dense, low-resolution panoramic images with encoded infrared light reflectivity and intensity shown in the bottom of Figure 1. This integration could further enhance the robustness and accuracy of odometry in complex environments. Although the current stage of fusion with LiDAR is still relatively early, it shows promising potential for future developments.
Acknowledgment
This research is supported by the Research Council of Finland’s Digital Waters (DIWA) flagship (Grant No. 359247) as well as the DIWA Doctoral Training Pilot project funded by the Ministry of Education and Culture (Finland).